OpenStudy (anonymous):

A difficult question can anyone find the length of the following curve? x^(2/3)+y^(2/3) =1

OpenStudy (anonymous):

need to do it manually.

OpenStudy (anonymous):

First solve for y. \[y ^{2/3}=1-x ^{2/3}\] Cube both sides \[y ^{2}=1-x ^{2}\] \[y=\sqrt{1-x ^{2}}\]

OpenStudy (anonymous):

Then use arc length formula \[\int\limits_{a}^{b}1+y ^{'2}\]

OpenStudy (anonymous):

good idea. But I'm afraid your cube is not correct. (1-x^(2/3))^3 doenst give you (1-x^(2))

OpenStudy (anonymous):

actually. we can do this by symmetry. find out one part of the curve first. yes. find out y and then dy/dx first. after that we can use the length formula.

OpenStudy (anonymous):

yea. \[y=\sqrt{1-x ^{2/3}}(1-x ^{2/3})\]

OpenStudy (anonymous):

did you get 3 as your final answer?

OpenStudy (anonymous):

What are your limits?

OpenStudy (anonymous):

0 to 1?

OpenStudy (anonymous):

bounded by 1 and 0. solve the length of the curve x belongs to (0,1) first.

OpenStudy (anonymous):

then times 2. symmetry

OpenStudy (anonymous):

Yea the integration gives you 3

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

no prob